OpenStudy (anonymous):

Solve for the given domain. Write the solution as a set of ordered pairs. 2x – 2y = 3, domain {-1, 0, 1, 2}. {(-1, -2.5), (0, -1.5), (1, .5), (2, .5)} {(-1, -2.5), (0, 1.5), (1, -.5), (2, .5)} {(-1, -2.5), (0, -1.5), (1, -.5), (2, .5)} {(-1, 2.5), (0, 1.5), (1, -2.5), (2, .5)}

OpenStudy (anonymous):

give that vakues for the eq Eg : {-1, 0, 1, 2}. select -1 2x - 2y =3 put x = -1 find y -2 - 2y = 3 -2y = 5 y = -5/2 \

OpenStudy (anonymous):

Do..the same thing for the rest